# Performance evaluation of compact plastic scintillating fiber modules for muon tomography applications

### v1

@@ Abstract
Muon tomography is a non-destructive imaging technique that exploits cosmic-ray muons from multiple directions. Its performance critically relies on stable, large-area, and high-resolution position-sensitive detectors. In this work, we report on the development of four compact scintillating fiber modules, each 100 cm long and composed of two staggered layers of 1 mm diameter fibers. The fibres are read out at one end by one-dimensional silicon photomultiplier arrays with a 2 mm pitch, coupled to Citiroc1A-based front-end electronics. The modules were characterised with cosmic-ray muons, yielding a detection efficiency above 97% and a mean spatial resolution of about 0.56 mm, with uniform response over different distances from the readout end. These results demonstrate the suitability of this detector design for compact and large-area systems in muon tomography applications.
@@ 1 Introduction
Cosmic-ray muons are secondary particles naturally produced by the interactions of primary cosmic rays with the Earth’s atmosphere. At sea level, their flux is typically about $\sim$$\rm{1\,cm^{-2}\,min^{-1}}$, with an average energy of $\sim$4 GeV [1]. Owing to their strong penetrating power, muon tomography can non-destructively reconstruct the internal structures of large and dense objects by tracking these muons from multiple directions [2, 3, 4]. Depending on the physical principle, the technique can be classified into muon absorption tomography, which exploits the muon transmission rate, and muon scattering tomography, which relies on the measurement of muon deflection angles. In recent years, muon tomography has been successfully applied in diverse fields, including archaeology [5], geological exploration [6], nuclear waste monitoring [7], and border security [8].
@@
Given the relatively low flux of cosmic-ray muons and the complexity of practical application scenarios, muon tomography systems demand detectors that combine high spatial resolution with large-area coverage and long-term operational stability. Plastic scintillator detectors are often preferred in such applications, as they can be flexibly manufactured in various geometries, maintain stable performance across different temperatures, offer high detection efficiency, and allow relatively low-cost large-area production. To achieve good spatial resolution, scintillators are usually cut into long bars with rectangular [9] or triangular [10, 11] cross sections and read out with wavelength-shifting fibers. New 3D printing techniques have also been developed to manufacture scintillators with more complex geometries [12]. However, the spatial resolution of such scintillator detectors is typically at the millimeter level, and further improvement toward the sub-millimeter scale is difficult due to the constraints imposed by the bar shape and lateral size.
@@
The plastic scintillating fiber (SciFi) detector offers much finer granularity compared to bulk scintillator detectors, enabling superior spatial resolution. In addition, its long attenuation length and high light yield make it possible to construct SciFi detectors with lengths of several meters. These features render the SciFi detector a promising candidate for compact muon tomography systems. This potential has been demonstrated in Refs. [13, 14] with round scintillating fibers, where grouped fibers were coupled with MAPMT or SiPM arrays to reduce the number of readout channels. To further enhance compactness and eliminate the laborious fiber-grouping process, we propose a SciFi detector employing a one-dimensional SiPM array readout [15] combined with a potential SiPM multiplexing strategy [16] for sub-millimeter spatial resolution. In this work, we focus on the design and performance evaluation of SciFi modules with SiPM array readout, while the multiplexing strategy will be discussed in a separate study.
@@
The remainder of this paper is organized as follows. Section 2 describes the design of the 100-cm-long SciFi modules aiming for sub-millimeter spatial resolution. Section 3 presents the experimental setup and data acquisition system used for performance evaluation. Section 4 details the data analysis procedures, including pedestal substraction, gain calibration, signal clustering, tracking, and alignment. Section 5 reports the experimental results on the detection efficiency and spatial resolution of the SciFi modules at different positions. Finally, Section 6 concludes the work and provides an outlook for future developments.
@@ 2 Construction of SciFi modules
The spatial resolution of SciFi detectors with one-dimensional SiPM array readout is generally determined by both the fibre diameter and the SiPM pitch. To reach sub-millimetre resolution, the preferred configuration is 1 mm diameter fibres coupled to SiPM arrays with a $1\times 2$ mm or $1\times 3$ mm size. However, due to the unavailability of such commercial SiPM products, the present SciFi modules are designed as shown in Fig. 1, employing two staggered layers of 1 mm diameter fibres read out by SiPM arrays with a $2\times 2$ mm size.
@@
Each compact SciFi module has dimensions of $5.5\times 100$ cm2 and is constructed using double-clad Kuraray SCSF-78M fibres [17]. These fibers have a typical trapping efficiency of about 5.4%, an attenuation length exceeding 4 m, and a light yield of roughly 8000 photons/MeV. Their emission spectrum spans 400–700 nm, peaking near 450 nm. To ensure precise positioning along the full module length, an aluminium alloy mould with precision-machined grooves is fabricated (Fig. 2(a)). After applying a release agent, the first fibre layer is placed in the grooves (Fig. 2(b)), followed by an adhesive mixed with $\rm TiO_{2}$ powder to reduce cross talks between fibers. The second layer is then positioned in the interstitial grooves. The assembly is covered with Teflon sheets, and demoulded once the adhesive solidifies. The fibre ends are cut and polished to ensure efficient optical coupling. Each module is mounted on a carbon-fibre support plate, with the readout end directly coupled to the SiPM arrays and the opposite end covered with enhanced specular reflector (ESR) film via an air gap (Fig. 2(c)). In addition, four SciFi modules of 10 cm length are produced using the same procedure but without ESR film to provide muon triggers in later experimental test.
To ensure efficient photon collection from the fibres, the dead area between adjacent SiPMs must be kept below 1 mm and minimized as much as possible. For this reason, Hamamatsu S13360-2050VE devices [18] with a $\rm 2\times 2$ mm2 photosensitive area were selected to assemble the one-dimensional arrays as shown in Fig. 2 (e). These devices employ through-silicon via (TSV) technology instead of conventional wire bonding, reducing the distance between the package edge and the photosensitive area to 0.2 mm on all four sides. An additional 0.1 mm margin was reserved to facilitate precise alignment and reliable soldering. The SiPMs are sensitive over the 320–900 nm wavelength range, matching well with the emission spectrum of the scintillating fibres. Their typical breakdown voltage is about 53 V and have peak photon detection efficiency reaches about 40% at 3 V overvoltage around 450 nm. Furthermore, each channel has a 50 $\rm\mu m$ pixel pitch and 1584 pixels, ensuring a linear response to photons emitted by the fibres when traversed by cosmic-ray muons.
@@ 3 Experimental setup
Figure 3: Experimental setup of the SciFi modules with Citiroc1A-based readout electronics. The four short modules provide the muon trigger for the long SciFi modules during cosmic-ray muon test. Their positions can be adjusted to record muon events at different locations along the long modules.
In the experimental setup shown in Fig. 3, four long SciFi modules of 100 cm length (layer 1-4), are arranged vertically with a spacing of approximately 5 cm along the $z$-axis and sensitive to $y$ positions of particles. Four shorter modules (layers 5–8), each 10 cm in length, are positioned perpendicularly to the long modules. Each short module can be slid along the $x$-axis to select muon trajectories at different positions along the long modules. The closest pair of short and long modules is separated by a vertical gap of about 0.5 mm and used to measure the two-dimensional position of muons. The long and short modules are read out by separate DT5550W data acquisition boards from CAEN, each equipped with four Citiroc 1A chips[19] on a piggyboard and a 14-bit ADC. The SiPM array of each SciFi module is connected to a dedicated Citiroc 1A chip and biased at 57 V via the piggyboard’s power supply module.
The Citiroc 1A is a 32-channel front-end ASIC optimized for SiPM readout. Each channel includes two parallel preamplifiers, one in high-gain (HG) mode and the other in low-gain (LG) mode, extending the dynamic range up to about 2500 photoelectrons (p.e.). Each preamplifier output is processed by a slow shaper with an adjustable shaping time between 12.5 ns and 87.5 ns, followed by a peak-sensing circuit to measure the integrated charge. For triggering, the chip provides a 15ns peaking-time fast shaper capable of discriminating signals down to 1/3 p.e. in each channel, as well as an ASIC-level charge trigger formed by the logical OR of all 32 channels. The DT5550W board can then generate a global trigger by combining the logic OR outputs of its four Citiroc 1A chips.
@@
  • Pedestal mode: SiPM signals are recorded without bias voltage, using 1 kHz periodic triggers generated by the DT5550W software.
  • Gain-calibration mode: The two DT5550W boards operates independently. A global trigger is issued whenever the charge in any channel of the four Citiroc 1A chips exceeds about 5 p.e., in which case all SiPM signals from either the four long or the four short modules were read out.
  • Muon-test mode: The DT5550W board for the short modules operates with ASIC-level charge trigger, issuing a global trigger whenever a channel of any Citiroc1A chip exceeded about 10 p.e. The threshold is chosen as a compromise between suppressing background events and retaining muon events. This trigger is sent to the DT5550W board of the long modules as an external trigger. Its trigger latency, including cables, is measured to be about 40 ns using charge injection into a Citiroc 1A under identical acquisition settings. This ensures that the trigger arrives before the peak-sensing stage of the long modules with 87.5 ns shaping time. Upon receiving the external trigger, the long-module data are read out. Complete events are finally built by matching long- and short-module data with the same trigger ID in offline analysis.
@@ 4.1 Channel calibration
Although in the cosmic-ray muon test all SiPM channels of a SciFi module were biased at the same voltage, variations in their breakdown voltages and the responses of different Citiroc1A channels resulted in non-uniform responses of number of photonelectrons ($N_{\mathrm{pe}}$) among SiPM channels. To account for these variations, the pedestal must be subtracted from the raw data and the gain of each SiPM channel calibrated individually.
@@
Figure 4: (Left) The raw ADC spectrum of one HG amplifier channel to extract gain and pedestal values. (Right) LG signal after pedestal substraction as a function of the calibrated HG $N_{\mathrm{pe}}$ for a representative SiPM channel, from which the LG gain is obtained by the linear fit.
As shown in the left panel of Fig. 4, the raw ADC spectrum of the HG amplifier signals obtained in gain-calibration mode exhibits a clear offset from zero. The pedestal is defined as the ADC value of the first and highest peak (from left to right), which is consistent with the value measured in pedestal mode with no high voltage applied to the SiPM arrays. The same procedure is applied for the pedestal estimation of the LG amplifier signals. The charge threshold set during gain-calibration mode corresponds to the valley at approximately 5 p.e. The peaks between the pedestal and the threshold are identified, and the HG gain, $G_{HG}$, is determined as the mean ADC separation between adjacent peaks. As reported in Refs. [20, 21], the ADC difference between the pedestal and the 1 p.e. peak is systematically smaller than the actual gain, since the peak-sensing stage always samples the maximum signal value within the hold-delay window. Consequently, the 1 p.e. peak can be treated as a “fake pedestal” to estimate the number of photoelectrons in an HG signal as
$N_{\mathrm{pe}}=\frac{ADC_{HG}-ADC_{1\mathrm{pe}}}{G_{HG}}+1,.$ (4.1)
The LG gain is intercalibrated with respect to the HG signals, as illustrated in the right panel of Fig. 4. It is obtained as the slope of a linear fit using HG signals in the range of 10–90 p.e. The ratio of obtained HG gain and LG gain as shown in Fig. 4 is consistent with that of amplifier value set in firmware. For larger SiPM signals (>90>90 p.e.), the HG response becomes non-linear, and in such cases the LG signals are used to estimate the input $\rm{N_{pe}}$. After this calibration, the raw ADC data acquired in muon-test mode are converted into $N_{\mathrm{pe}}$ values. As shown in the left panel of Fig. 5, all 22 SiPM channels of layer 1 exhibit consistent signal spectra after pedestal subtraction and gain calibration. Moreover, the distribution of the SiPM channel with the maximum $N_{\mathrm{pe}}$ in each event is consistent across all four long modules, as illustrated in the right panel of Fig. 5. The slight deviation observed in layer 2 is attributed to differences in the procedure used when attaching the ESR films.
@@
Figure 5: (Left) Calibrated SiPM signals of layer 1 acquired in muon-test mode after pedestal subtraction and gain calibration. (Right) Distribution of the SiPM channel with the maximum $N_{\mathrm{pe}}$ in each event for all four long modules when the short modules were centered at 52.5 cm from the readout end of the long modules.

### v2

@@ Abstract
Muon tomography is a non-destructive imaging technique that exploits cosmic-ray muons from multiple directions. Its performance critically depends on the stability, active-area coverage, and spatial resolution of position-sensitive detectors. In this work, we report on the development of four compact scintillating fiber modules, each 100 cm long and composed of two staggered layers of 1 mm diameter fibers. The fibers are read out at one end by one-dimensional silicon photomultiplier arrays with a 2 mm pitch, coupled to Citiroc 1A-based front-end electronics. The modules were characterized with cosmic-ray muons, yielding a detection efficiency above 97% and a mean spatial resolution of about 0.56 mm, with uniform response over different distances from the readout end. An imaging test of a lead block was also performed, and the reconstructed results are consistent with the block’s profile. These results demonstrate the suitability of this detector design for compact and large-area systems in muon tomography applications.
keywords:
plastic scintillating fiber detector, detection efficiency, position resolution, muon tomography
[…]
@@ 1 Introduction
Cosmic-ray muons are secondary particles naturally produced by the interactions of primary cosmic rays with the Earth’s atmosphere. At sea level, their flux is typically $\sim$$\rm{1\,cm^{-2}\,min^{-1}}$, with an average energy of $\sim$4 GeV [Navas:2024]. Owing to their strong penetrating power, Owing to their strong penetrating power, cosmic-ray muons enable muon tomography to non-destructively reconstruct the internal structures of large and dense objects by tracking these muons from multiple directions [Borozdin2003, Bonomi:2020, Tanaka:2023]. Depending on the physical principle, the technique can be classified into muon absorption tomography, which exploits the muon transmission rate, and muon scattering tomography, which relies on the measurement of muon deflection angles. In recent years, muon tomography has been successfully applied in diverse fields, including archaeology [Morishima:2017ghw], geological exploration [Liu:2024], nuclear waste monitoring [Mahon:2018], and border security [Barnes:2023].
@@
Because of the relatively low flux of cosmic-ray muons and the complexity of practical environments, muon tomography systems demand detectors that combine high spatial resolution with large-area coverage and long-term operational stability. Plastic scintillator detectors are often preferred in such applications, as they can be flexibly manufactured in various geometries, maintain stable performance across different temperatures, offer high detection efficiency, and allow relatively low-cost large-area production. To achieve good spatial resolution, scintillators are usually cut into long bars with rectangular [Lesparre:2012] or triangular [Anstasio:2013, Wang:2024] cross sections, which are read out with wavelength-shifting fibers. New 3D printing techniques have also been developed to manufacture scintillators with more complex geometries [Yu:2025]. However, the spatial resolution of such scintillator detectors is typically at the millimeter level, and further improvement toward the sub-millimeter scale is difficult due to the constraints imposed by the bar shape and lateral size.
@@
The plastic scintillating fiber (SciFi) detector offers much finer granularity compared to bulk scintillator detectors, enabling superior spatial resolution. In addition, its long attenuation length and high light yield make it possible to construct SciFi detectors with lengths of several meters. These features render the SciFi detector a promising candidate for compact muon tomography systems. This potential has been demonstrated in Refs. [Clarkson:2014, Anbarjafari:2021] with round scintillating fibers, where grouped fibers were coupled with multi-anode PMTs or SiPM arrays to reduce the number of readout channels. To further enhance compactness and eliminate the laborious fiber-grouping process, we propose a SciFi detector employing a one-dimensional SiPM array readout [Chen:2023] combined with a potential SiPM multiplexing strategy [lv2025multiplexed] for sub-millimeter spatial resolution. In this work, we focus on the design and performance evaluation of SciFi modules with SiPM array readout, while the multiplexing strategy is discussed in a separate study.
@@
The remainder of this paper is organized as follows. Section 2 describes the design of the 100-cm-long SciFi modules aiming for sub-millimeter spatial resolution. Section 3 presents the experimental setup and data acquisition system used for performance evaluation. Section 4 details the data analysis procedures, including pedestal subtraction, gain calibration, signal clustering, tracking, and alignment. Section 5 reports the experimental results on the detection efficiency and spatial resolution of the SciFi modules at different positions. Section 6 demonstrates a muon scattering tomography test and evaluates the imaging performance of the modules. Finally, Section 7 concludes the work and provides an outlook for future developments.
@@ 2 Construction of SciFi modules
The spatial resolution of SciFi detectors with one-dimensional SiPM array readout is primarily determined by both the fiber diameter and the SiPM pitch. To achieve sub-millimetre precision, the preferred configuration would employ 1 mm diameter fibers coupled to SiPM arrays with a $1\times 2$ $\rm{mm^{2}}$ or $1\times 3$ $\rm{mm^{2}}$ active area per channel. However, such SiPM products are not yet commercially available. Therefore, as illustrated in Fig. 1, the present SciFi modules are designed with two staggered layers of 1 mm diameter fibers, read out by SiPM arrays with a $2\times 2$ $\rm{mm^{2}}$ active area per channel.
@@
Each compact SciFi module has dimensions of 100 cm (length) $\times$ 5.5 cm (width) $\times$ 2 mm (thickness) and is constructed using double-clad Kuraray SCSF-78M fibers [kurarayScintillatingFibers]. These fibers have a typical trapping efficiency of about 5.4%, an attenuation length exceeding 4 m, and a light yield of roughly 8000 photons/MeV. Their emission spectrum spans 400–700 nm, peaking near 450 nm. To ensure a precise positioning throughout the module length, an aluminum alloy mold with precision-machined grooves is fabricated (Fig. 2(a)). After applying a release agent, the first fiber layer is placed in the grooves (Fig. 2(b)), followed by an adhesive mixed with $\rm TiO_{2}$ powder to reduce optical crosstalk between fibers. The second layer is then positioned in the interstitial grooves. The assembly is covered with Teflon sheets and is demolished once the adhesive solidifies. The fiber ends are cut and polished to ensure efficient optical coupling. Each module is mounted on a carbon-fiber support plate, with the readout end directly coupled to the SiPM arrays and the opposite end covered with an enhanced specular reflector (ESR) film via an air gap (Fig. 2(c)). In addition, four short SciFi modules (10 cm (length) $\times$ 5.5 cm (width) $\times$ 2 mm (thickness)) are also produced using the same procedure but without ESR film to serve as muon trigger detectors in later tests.
To ensure efficient photon collection from the fibers, the dead area between adjacent SiPMs must be kept below 1 mm and minimized as much as possible. For this reason, Hamamatsu S13360-2050VE devices [HamamatsuS13360-2050VE] with a $\rm{2\times 2}$ $\rm{mm^{2}}$ photosensitive area are selected to assemble the one-dimensional arrays as shown in Fig. 2(e). These devices employ through-silicon via (TSV) technology instead of conventional wire bonding, reducing the distance between the package edge and the photosensitive area to 0.2 mm on all four sides. An additional 0.1 mm margin was reserved to facilitate precise alignment and reliable soldering. The SiPMs are sensitive over the 320–900 nm wavelength range, matching well with the emission spectrum of the scintillating fibers. Their typical breakdown voltage is about 53 V and have peak photon detection efficiency reaches about 40% at 3 V overvoltage around 450 nm. Furthermore, each channel has 50 $\rm\mu m$ pixel pitch and 1584 pixels, ensuring a linear response to photons emitted by the fibers when traversed by cosmic-ray muons.
@@ 3 Experimental setup
Figure 3: Experimental setup of the SciFi modules with Citiroc1A-based readout electronics. The four short modules provide the muon trigger for the long SciFi modules during the cosmic-ray muon test. Their positions can be adjusted to record muon events at different locations along the long modules.
The experimental setup is shown in Fig. 3. Four long SciFi modules (layers 1–4), each 100 cm in length and aligned along the $x$-axis, are arranged vertically with a spacing of approximately 5 cm along the $z$-axis. These layers are sensitive to the $y$-coordinate of incident particles. Four shorter modules (layers 5–8), each 10 cm in length and oriented along the $y$-axis, are placed perpendicularly to the long modules. The short modules can be translated as a group along the $x$-axis to scan muons striking different positions along the length of the long modules. The closest pair of short and long modules is separated by a vertical gap of about 0.5 mm, allowing for two-dimensional position measurements of muons. The long and short modules are read out by separate DT5550W data acquisition boards from CAEN, each equipped with four Citiroc 1A chips[Citiroc1A] on a piggyback board and a 14-bit ADC. The SiPM array of each SciFi module is connected to a dedicated Citiroc 1A chip and biased at 57 V via the piggyback’s power supply module.
The Citiroc 1A is a 32-channel front-end ASIC optimized to amplify and integrate SiPM signals. Each channel includes two parallel preamplifiers, one in high-gain (HG) mode and the other in low-gain (LG) mode, extending the dynamic range up to about 2500 photoelectrons (p.e.). Each preamplifier output is processed by a slow shaper with an adjustable shaping time between 12.5 ns and 87.5 ns, followed by a peak-sensing circuit to measure the integrated charge. For triggering, the chip provides a 15 ns peaking-time fast shaper capable of discriminating signals down to 1/3 the signal amplitude of one p.e. in each channel, as well as an ASIC-level charge trigger formed by the logical OR of all 32 channels. The DT5550W board can then generate a global trigger by combining the logic OR outputs of its four Citiroc 1A chips.
@@
  • 1.
    Pedestal mode: SiPM signals were recorded without bias voltage using 1 kHz periodic triggers generated by the DT5550W software.
  • 2.
    Gain-calibration mode: In this mode, the two DT5550W boards operated independently. A global trigger was issued whenever the charge in any channel of the four Citiroc 1A chips exceeded about 5 p.e., prompting the readout of all SiPM channels of the four long (or short) modules by their respective DAQ boards. The threshold of 5 p.e. was empirically chosen to maintain a manageable trigger rate. Under these conditions, most recorded signals originated from SiPM dark noise or ambient radiation rather than cosmic-ray muons.
  • 3.
    Muon-test mode: For the cosmic-ray muon tests, the DT5550W board connected to the short modules was configured to generate ASIC-level charge triggers whenever any Citiroc 1A channel exceeded about 10 p.e. This threshold was chosen as a compromise between background suppression and muon detection efficiency. The trigger signal was then sent to the DT5550W board of the long modules as an external trigger. The total trigger latency, including cable delays, was measured to be approximately 40 ns using charge injection into a Citiroc 1A under identical acquisition conditions. This ensured that the trigger arrived prior to the peak-sensing stage of the long modules with 87.5 ns shaping time. Upon receiving the external trigger, the long-module data were read out. Complete events were subsequently constructed offline by matching the long- and short-module data according to their trigger IDs. Figure 4 shows the hourly trigger rate of the experimental setup operated in muon-test mode, demonstrating stable and reliable performance over the full data-taking period.
    Figure 4: Hourly trigger rate of the experimental setup operated in muon-test mode, demonstrating stable and reliable performance.
@@ 4.1 Channel calibration
The raw signal from a fired SiPM is the sum of signals from all microcells that fired in Geiger mode. The number of fired microcells corresponds to the number of photoelectrons ($N_{\rm{pe}}$) for the channel, making $N_{\rm{pe}}$ proportional to the integral of the SiPM pulse, i.e., the total charge. Although all SiPM channels of a SciFi module were biased at the same voltage during the cosmic-ray muon test, variations in their breakdown voltages and the responses of different Citiroc1A channels resulted in non-uniform responses in $N_{\mathrm{pe}}$ among SiPM channels. To account for these variations, the pedestal must be subtracted from the raw data, and the gain of each SiPM channel must be calibrated individually.
@@
Figure 5: (Left) The raw ADC spectrum of one HG amplifier channel used to extract gain and pedestal values. (Right) LG signal after pedestal subtraction as a function of the calibrated HG $N_{\mathrm{pe}}$ for a representative SiPM channel, from which the gain in LG mode is obtained by a linear fit.
As shown in the left panel of Fig. 5, the raw ADC spectrum of the HG amplifier signals obtained during gain calibration clearly exhibits an offset from zero. The pedestal is defined as the ADC value of the first and highest peak (from left to right), which is consistent with the value measured in pedestal mode with no high voltage applied to the SiPM arrays. The same procedure is applied to estimate the pedestal for the LG amplifier signals. The charge threshold set during gain calibration corresponds to the valley at approximately 5 p.e. The peaks between the pedestal and the threshold are identified, and the high gain, $G_{HG}$, is determined as the mean ADC separation between the adjacent peaks. As reported in Refs. [Impiombato:2020, Wu:2024], the ADC difference between the pedestal and the 1 p.e. peak is systematically smaller than the actual gain, since the peak-sensing stage always samples the maximum signal value within the hold-delay window. Consequently, the 1 p.e. peak can be treated as a “fake pedestal” to estimate $N_{\rm{pe}}$ in an HG signal, as given by the equation:
$N_{\mathrm{pe}}=\frac{ADC_{HG}-ADC_{1\mathrm{pe}}}{G_{HG}}+1.$ (1)
The gain in LG mode is inter-calibrated relative to the HG signals, as illustrated in the right panel of Fig. 5. It is obtained as the slope of a linear fit using HG signals in the range of 10–90 p.e. The ratio of the obtained gains in HG and LG modes, as shown in Fig. 5, is consistent with the amplifier value set in firmware. For larger SiPM signals (>90>90 p.e.), the HG response becomes non-linear. In such cases, LG signals are used to estimate the input $\rm{N_{pe}}$. After this calibration, the raw ADC data acquired in muon-test mode are converted into $N_{\mathrm{pe}}$ values. As shown in the left panel of Fig. 6, all 22 SiPM channels of layer 1 exhibit consistent signal spectra after pedestal subtraction and gain calibration. Moreover, the distribution of the SiPM channel with the maximum $N_{\mathrm{pe}}$ in each event is consistent across all four long modules, as illustrated in the right panel of Fig. 6. The slight deviation observed in layer 2 is attributed to differences in the procedure used when attaching the ESR films.
@@
Figure 6: (Left) Calibrated SiPM signals from layer 1 acquired in muon-test mode after pedestal subtraction and gain calibration. (Right) Distribution of the SiPM channel with the maximum $N_{\mathrm{pe}}$ in each event for all four long modules when the short modules were centered at 52.5 cm from the readout end of the long modules.